# arXiv:2603.00917v1[cs.CL] 01 Mar 2026

### v1

@@ Abstract
Background: Artificial intelligence is increasingly deployed in healthcare workflows, and small open-source language models are gaining attention as viable tools for low-resource settings where cloud infrastructure is unavailable. Despite their growing accessibility, the reliability of these models — particularly the stability of their outputs under different phrasings of the same clinical question — remains poorly understood.
@@
Methods: We evaluated five open-source models — Gemma 2 (2B), Phi-3 Mini (3.8B), Llama 3.2 (3B), Mistral 7B, and Meditron-7B (domain-pretrained, not instruction-tuned) — across three clinical question answering datasets (MedQA, MedMCQA, and PubMedQA), sampling 200 questions from each. Each question was presented in five prompt styles (original, formal, simplified, roleplay, and direct). We measured a majority-based consistency score, overall and per-style accuracy, and instruction-following failure rate (UNKNOWN rate). All inference was conducted locally on consumer CPU hardware without domain-specific fine-tuning.
@@
Conclusions: High consistency does not imply correctness in small clinical language models — models can be reliably wrong, representing a particularly dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
@@ 1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, from clinical decision support to exam-style medical question answering. While much attention has focused on large, cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). However, as these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but on how stable and predictable their outputs are under different conditions. Despite rapid adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
@@
Most existing evaluations of clinical language models focus primarily on accuracy — whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction matters: a model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
@@
To our knowledge, no prior study has systematically examined whether small, CPU-runnable models provide consistent answers under controlled prompt variation in clinical settings. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Recent surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs without analysing answer consistency as a reliability dimension. This study addresses that gap directly.
@@
In this work, we systematically evaluate prompt sensitivity in five open-source language models — Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained) — across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA) using five controlled prompt styles designed to simulate realistic query reformulation. We introduce a quantitative consistency score that measures how often a model produces the same answer across prompt variations for the same question, and analyse this alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare settings.
@@
We report four central findings. First, consistency and accuracy are independent metrics: models that produce highly stable answers across prompt variations are not necessarily more correct, with Gemma 2 demonstrating the clearest case of high consistency alongside low accuracy. Second, roleplay prompt styles systematically reduce performance across all models and datasets, indicating that persona-based framing negatively affects clinical question answering reliability. Third, instruction-following failure rates vary independently of model size, demonstrating that larger parameter counts do not guarantee safer or more dependable outputs under constrained deployment conditions. Fourth, Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction-following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical deployment.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on clinical LLM evaluation, prompt sensitivity, and small language models in healthcare. Section 3 describes our datasets, models, prompt variation design, inference setup, and evaluation metrics. Section 4 presents experimental results. Section 5 discusses clinical implications, limitations, and future directions. Section 6 concludes the paper.
@@ 2 Related Work
Most prior work evaluating large language models in healthcare has focused on accuracy, reasoning quality, and output safety. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through blinded physician preference on tasks including summarisation and question answering, demonstrating that smaller domain-specific models can outperform larger general-purpose systems on clinically relevant criteria. Kim et al. (Kim and others, 2025) systematically examine medical hallucinations in foundation models, highlighting the dangers of confident but incorrect outputs and their potential to mislead clinical decision-making. While these studies provide critical insights into output quality and safety, neither investigates whether models produce consistent answers when the same clinical question is phrased differently — leaving prompt sensitivity largely unexplored in healthcare AI.
Growing interest in small, CPU-runnable language models has emerged in response to the practical constraints of healthcare deployment in low-resource settings. Garg et al. (Garg and others, 2024) survey small language models in healthcare, documenting their efficiency advantages and suitability for local deployment under limited infrastructure. The MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) benchmarks provide standardised evaluation tasks across diverse clinical question formats, and have been widely adopted for assessing compact models. These works collectively advance the case for small model deployment in accessible healthcare AI, but focus primarily on task accuracy and computational efficiency rather than answer stability under prompt variation.
Prompt sensitivity — the tendency of model outputs to change in response to semantically equivalent but stylistically different inputs — has been studied in general NLP research. Ngweta et al. (Ngweta and others, 2024) demonstrate that minor changes in prompt formatting cause measurable performance fluctuations across multiple models and benchmarks, and propose mitigation strategies to reduce this brittleness. However, such investigations have not been extended to clinical question answering, and the safety implications of prompt sensitivity for healthcare AI deployment remain unexamined.
Taken together, these lines of research motivate the present study, which systematically evaluates prompt sensitivity and answer consistency in small open-source clinical language models — complementing prior work on accuracy, hallucination, and general prompt robustness with a focus on reliability under realistic deployment constraints.
@@ 3.1 Datasets
We evaluated model consistency and accuracy using three established clinical question answering benchmarks. MedQA (Jin et al., 2021) contains multiple-choice questions derived from the United States Medical Licensing Examination (USMLE), representing clinical knowledge required for medical licensure. MedMCQA (Pal et al., 2022) consists of multiple-choice questions from Indian medical entrance examinations (AIIMS/NEET), covering diverse medical specialties including pharmacology, anatomy, and pathology. Although the dataset includes subject labels, our evaluation used a randomly sampled subset of 200 questions without stratification by specialty. PubMedQA (Jin et al., 2019) presents biomedical research questions paired with PubMed abstracts, requiring yes, no, or maybe answers based on provided evidence rather than memorized knowledge. From each dataset we randomly sampled 200 questions using a fixed random seed (seed=42) to ensure reproducibility, yielding 600 questions total across three distinct clinical reasoning contexts.
@@ 3.2 Models
We evaluated five open-source small language models spanning a parameter range of 2B to 7B, including one domain-adapted medical pretraining model to examine the effect of model scale on prompt consistency. The selected models were Phi-3 Mini (Abdin and others, 2024) (3.8B parameters, Microsoft), Llama 3.2 (Grattafiori and others, 2024) (3B parameters, Meta), Gemma 2 (Gemma Team and others, 2024) (2B parameters, Google), and Mistral (Jiang and others, 2023) (7B parameters, Mistral AI). These models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families — enabling cross-architecture comparison of consistency behavior. All inference was conducted locally using Ollama without any domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behavior rather than task-specific adaptation.
In addition to these instruction-tuned general models, we included Meditron-7B (Chen and others, 2023), a domain-adapted language model developed by EPFL. Meditron-7B is built on Llama-2 and continued-pretrained on a curated medical corpus comprising PubMed articles, biomedical abstracts, and internationally recognised clinical practice guidelines. Critically, Meditron-7B is a base pretraining model without instruction tuning, meaning it has not been trained to follow structured task instructions. We include it explicitly to examine whether domain-specific medical pretraining alone — in the absence of instruction tuning — is sufficient for structured clinical question answering, and to contrast its behaviour with that of general instruction-tuned models. This comparison addresses a practical question for clinical AI deployment: is domain knowledge alone sufficient, or is instruction-following capability equally essential?
@@ 3.3 Prompt Variation Design
To systematically evaluate how small language models respond to differently worded clinical prompts, we designed a prompt variation engine that generates five semantically equivalent but stylistically distinct prompt formulations for each question.
@@
Each question across all three datasets was transformed into these five prompt styles, yielding 3,000 total prompts (600 questions $\times$ 5 styles). By keeping semantic content identical across variations, any differences in model output can be attributed to prompt sensitivity rather than question content.
@@ 3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures fully deterministic outputs — the model produces identical responses for identical prompts — which is essential for isolating prompt sensitivity as the source of answer variation rather than stochastic sampling. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures. All responses — including UNKNOWN outputs — were retained for consistency scoring and analysis.
@@ 3.5 Evaluation Metrics
We defined four complementary metrics to quantify model behavior under prompt variation.
Consistency Score. For each question, we identify the majority answer — the response selected most frequently across the five prompt styles — and compute the proportion of styles that agree with it. Formally, given responses $r_{1},r_{2},r_{3},r_{4},r_{5}$ for a question, the consistency score is:
@@
where $\hat{r}$ is the majority answer and $\mathbf{1}[\cdot]$ is the indicator function. A score of 1.0 indicates perfect consistency across all prompt styles; a score of 0.2 indicates maximum inconsistency. For example, responses [A, B, B, B, A] yield a majority answer of B and a consistency score of 3/5 = 0.60.
Overall Accuracy. The majority answer for each question is compared against the dataset ground-truth label, reflecting task-level correctness while accounting for variability across prompt styles.
@@
Together these metrics provide a holistic assessment of model behavior: a model may be highly consistent but consistently wrong, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
@@ 4.1 Overall Consistency and Accuracy
[…]
Statistical significance of observed differences was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. Consistency differences between models were statistically significant on MedQA and MedMCQA (p<<0.001 for most pairwise comparisons) but not on PubMedQA, suggesting that dataset characteristics influence the degree of prompt sensitivity. Accuracy differences were generally non-significant on MedQA, while on PubMedQA, Llama 3.2 significantly outperformed both Phi-3 Mini and Gemma 2 (p<<0.001), confirming its superior clinical reasoning on biomedical evidence-based questions. Full significance test results are provided in Table LABEL:tab:stats.
@@
MedQA Llama 3.2 Phi-3 Cons. 0.0005 ***
Gemma 2 Phi-3 Cons. <<.0001 ***
Mistral Phi-3 Cons. <<.0001 ***
@@
Llama 3.2 Phi-3 Acc. 0.9020 ns
Gemma 2 Phi-3 Acc. 0.0689 ns
Mistral Phi-3 Acc. 0.5383 ns
@@
MedMCQA Llama 3.2 Phi-3 Cons. 0.0437 *
Gemma 2 Phi-3 Cons. <<.0001 ***
Mistral Phi-3 Cons. 0.0001 ***
@@
Llama 3.2 Phi-3 Acc. 0.5962 ns
Gemma 2 Phi-3 Acc. 0.0327 *
Mistral Phi-3 Acc. 0.1060 ns
@@
PubMedQA Llama 3.2 Phi-3 Cons. 0.1055 ns
Gemma 2 Phi-3 Cons. 0.8497 ns
Mistral Phi-3 Cons. 0.6172 ns
@@
Llama 3.2 Phi-3 Acc. <<.0001 ***
Gemma 2 Phi-3 Acc. 0.0002 ***
Mistral Phi-3 Acc. 0.1930 ns
@@
Accuracy patterns diverged substantially from consistency rankings (Figure 2). Llama 3.2 achieved the highest overall accuracy across datasets (49.0–65.0%), followed by Phi-3 Mini (48.0–53.0%) and Mistral 7B (42.5–45.5%). Gemma 2, despite exhibiting the highest consistency scores, achieved the lowest accuracy (33.0–43.5%). This inverse relationship between consistency and accuracy in Gemma 2 highlights a critical finding: high consistency does not imply high accuracy. Models can be reliably wrong — producing the same incorrect answer across all prompt variations — which represents a particularly dangerous failure mode in clinical decision support contexts.
Meditron-7B, included as a domain-adapted pretraining model without instruction tuning, exhibited markedly different behaviour from all instruction-tuned models. While it achieved moderate consistency on MedQA and MedMCQA (0.761 and 0.762 respectively), its UNKNOWN rates were substantially higher (22.8% and 22.4%), indicating frequent instruction-following failures. On PubMedQA, Meditron-7B produced near-complete failure — a consistency score of 0.010, 2.0% accuracy, and 99.0% UNKNOWN rate — suggesting the model is unable to produce structured yes/no/maybe responses without instruction tuning. These results indicate that domain-specific pretraining alone is insufficient for structured clinical question answering, and that instruction-following capability is a prerequisite for clinical deployment.

### v2

# Prompt Sensitivity and Answer Consistency of Small Open-Source Large Language Models on Clinical Question Answering: Implications for Low-Resource Healthcare Deployment

@@ Abstract
Background: Artificial intelligence is increasingly deployed in healthcare workflows, and small open-source language models are gaining attention as viable tools for low-resource settings where cloud infrastructure is unavailable. Despite their growing accessibility, the reliability of these models, particularly the stability of their outputs under different phrasings of the same clinical question, remains poorly understood.
@@
Methods: Five open-source language models were evaluated across three clinical question answering datasets (MedQA, MedMCQA, PubMedQA) using five prompt styles. Consistency scores, accuracy, and instruction-following failure rates were measured using experiments conducted locally on consumer CPU hardware without fine-tuning.
@@
Conclusions: High consistency does not imply correctness in small clinical language models; models can be reliably incorrect, representing a particularly dangerous failure mode in clinical decision support. Roleplay prompt styles should be avoided in healthcare AI applications. Among the models evaluated, Llama 3.2 demonstrated the strongest balance of accuracy and reliability for low-resource deployment. These findings highlight the necessity of multidimensional evaluation frameworks that assess consistency, accuracy, and instruction adherence jointly for safe clinical AI deployment.
@@ 1 Introduction
Artificial intelligence is increasingly being integrated into healthcare workflows, from clinical decision support to exam-style medical question answering. While much attention has focused on large, cloud-based models, smaller open-source language models are gaining importance because they can run locally on standard CPUs, making them viable for rural clinics, community hospitals, and low-resource health systems with limited infrastructure (Garg and others, 2024). However, as these smaller models become more accessible for real-world deployment, questions about their reliability become critical. In clinical settings, even small inaccuracies can have significant consequences, and the safety of AI systems depends not only on their overall performance but also on how stable and predictable their outputs are under different conditions. Despite rapid adoption, the reliability of small, locally deployable models in healthcare remains insufficiently studied.
@@
Most existing evaluations of clinical language models focus primarily on accuracy, that is, whether a model selects the correct answer on a given benchmark (Bedi and others, 2025). However, accuracy alone does not capture how stable a model’s outputs are when the same clinical question is phrased differently. In safety-critical domains such as healthcare, this distinction matters. A model that gives different answers depending on how a question is worded cannot be considered reliable, regardless of its aggregate accuracy score.
@@
To our knowledge, no prior study has systematically evaluated whether small, CPU-runnable language models provide consistent answers under controlled prompt variation in clinical settings. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through physician preference but does not assess response stability across prompt reformulations. Ngweta et al. (Ngweta and others, 2024) quantify prompt brittleness in general NLP tasks but do not examine clinical datasets or safety implications. Kim et al. (Kim and others, 2025) demonstrate that foundation models frequently produce confident but incorrect medical outputs, yet do not empirically measure variability under prompt changes. Recent surveys on small language models in healthcare (Garg and others, 2024) discuss efficiency and deployment tradeoffs without analyzing answer consistency as a reliability dimension. The present study addresses this gap directly.
@@
The present work systematically evaluates prompt sensitivity in five open-source language models, including Phi-3 Mini (3.8B), Llama 3.2 (3B), Gemma 2 (2B), Mistral 7B, and Meditron-7B (domain-pretrained), across three established clinical question answering benchmarks (MedQA, MedMCQA, and PubMedQA) using five controlled prompt styles designed to simulate realistic query reformulation. A quantitative consistency score is introduced that measures how often a model produces the same answer across prompt variations for the same question and analyzes this alongside overall accuracy, per-style accuracy, and instruction-following failure rates. All experiments were conducted locally on consumer CPU hardware using unmodified base models without medical fine-tuning, reflecting realistic deployment conditions in low-resource healthcare settings.
@@
Four central findings emerge from the analysis. First, consistency and accuracy are independent metrics: models that produce highly stable answers across prompt variations are not necessarily more correct, with Gemma 2 demonstrating the clearest case of high consistency alongside low accuracy. Second, roleplay prompt styles systematically reduce performance across all models and datasets, indicating that persona-based framing negatively affects clinical question answering reliability. Third, instruction-following failure rates vary independently of model size, demonstrating that larger parameter counts do not guarantee safer or more dependable outputs under constrained deployment conditions. Fourth, Meditron-7B, a domain-pretrained model without instruction tuning, exhibited near-complete instruction- following failure despite encoding substantial medical knowledge, demonstrating that domain knowledge and instruction-following capability are orthogonal requirements for clinical deployment.
@@
The remainder of this paper is structured as follows. Section 2 reviews related work on clinical LLM evaluation, prompt sensitivity, and small language models in healthcare. Section 3 describes the datasets, models, prompt variation design, inference setup, and evaluation metrics. Section 4 presents experimental results. Section 5 discusses clinical implications, limitations, and future directions. Section 6 concludes the paper.
@@ 2 Related Work
Most prior work evaluating large language models in healthcare has focused on accuracy, reasoning quality, and output safety. The CLEVER framework (Kocaman and others, 2025) evaluates clinical LLM outputs through blinded physician preference on tasks including summarization and question answering, demonstrating that smaller domain-specific models can outperform larger general-purpose systems on clinically relevant criteria. Kim et al. (Kim and others, 2025) systematically examine medical hallucinations in foundation models, highlighting the dangers of confident but incorrect outputs and their potential to mislead clinical decision-making. While these studies provide critical insights into output quality and safety, neither investigates whether models produce consistent answers when the same clinical question is phrased differently, leaving prompt sensitivity largely unexplored in healthcare AI.
Growing interest in small, CPU-runnable language models has emerged in response to the practical constraints of healthcare deployment in low-resource settings. Garg et al. (Garg and others, 2024) survey small language models in healthcare, documenting their efficiency advantages and suitability for local deployment under limited infrastructure. The MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and PubMedQA (Jin et al., 2019) benchmarks provide standardized evaluation tasks across diverse clinical question formats and have been widely adopted for assessing compact models. These works collectively advance the case for small model deployment in accessible healthcare AI but focus primarily on task accuracy and computational efficiency rather than answer stability under prompt variation.
Prompt sensitivity, defined as the tendency of model outputs to change in response to semantically equivalent but stylistically different inputs, has been studied in general NLP research. Ngweta et al. (Ngweta and others, 2024) demonstrate that minor changes in prompt formatting cause measurable performance fluctuations across multiple models and benchmarks and propose mitigation strategies to reduce this brittleness. However, such investigations have not been extended to clinical question answering, and the safety implications of prompt sensitivity for healthcare AI deployment remain largely unexplored.
Taken together, these lines of research motivate the present study, which systematically evaluates prompt sensitivity and answer consistency in small open-source clinical language models. This work complements prior studies on accuracy, hallucination, and general prompt robustness by focusing on reliability under realistic deployment constraints.
@@ 3.1 Datasets
Model consistency and accuracy were evaluated using three established clinical question answering benchmarks. MedQA (Jin et al., 2021) contains multiple-choice questions derived from the United States Medical Licensing Examination (USMLE), representing clinical knowledge required for medical licensure. MedMCQA (Pal et al., 2022) consists of multiple-choice questions from Indian medical entrance examinations (AIIMS/NEET), covering diverse medical specialties including pharmacology, anatomy, and pathology. Although the dataset includes subject labels, the evaluation was conducted on a randomly sampled subset of 200 questions without stratification by specialty. PubMedQA (Jin et al., 2019) presents biomedical research questions paired with PubMed abstracts, requiring yes, no, or maybe answers based on provided evidence rather than memorized knowledge.
From each dataset, 200 questions were randomly sampled using a fixed random seed (seed = 42) to ensure reproducibility, yielding 600 questions total across three distinct clinical reasoning contexts.
@@ 3.2 Models
Five open-source language models spanning a parameter range of 2B to 7B parameters were evaluated, including one domain-adapted medical pretraining model to examine the effect of model scale on prompt consistency. The selected models were Phi-3 Mini (Abdin and others, 2024) (3.8B parameters, Microsoft), Llama 3.2 (Grattafiori and others, 2024) (3B parameters, Meta), Gemma 2 (Gemma Team and others, 2024) (2B parameters, Google), and Mistral (Jiang and others, 2023) (7B parameters, Mistral AI). These models were selected because they are fully open-source, capable of running on consumer CPU hardware without GPU acceleration, and represent distinct model families, enabling cross-architecture comparison of consistency behavior. All inference was conducted locally using Ollama without any domain-specific fine-tuning, ensuring that observed consistency patterns reflect inherent model behavior rather than task-specific adaptation.
In addition to these instruction-tuned general models, Meditron-7B (Chen and others, 2023), a domain-adapted language model developed by EPFL, was included. Meditron-7B is built on Llama-2 and further pretrained on a curated medical corpus comprising PubMed articles, biomedical abstracts, and internationally recognized clinical practice guidelines. Critically, Meditron-7B is a base pretraining model without instruction tuning, meaning it has not been trained to follow structured task instructions. It was explicitly included to examine whether domain-specific medical pretraining alone, in the absence of instruction tuning, is sufficient for structured clinical question answering and to contrast its behavior with that of general instruction-tuned models. This comparison addresses a practical question for clinical AI deployment: whether domain knowledge alone is sufficient or whether instruction-following capability is equally essential.
@@ 3.3 Prompt Variation Design
To systematically evaluate how small language models respond to differently worded clinical prompts, a prompt variation engine was designed to generate five semantically equivalent but stylistically distinct prompt formulations for each question.
@@
Each question across all three datasets was transformed into these five prompt styles, yielding 3,000 total prompts (600 questions $\times$ 5 styles). By keeping semantic content identical across variations, any differences in model output can be attributed to prompt sensitivity rather than differences in question content.
@@ 3.4 Inference Setup
All models were served locally via Ollama and queried through its REST API. Inference was conducted with temperature set to 0 and a maximum token limit of 10 tokens per response. Setting temperature to 0 ensures fully deterministic outputs, meaning the model produces identical responses for identical prompts. This is essential for isolating prompt sensitivity as the source of answer variation rather than stochastic sampling. Limiting output to 10 tokens prevents verbose responses and constrains the model to produce the single-letter (A, B, C, or D) or keyword (yes, no, or maybe) answer required by each dataset format.
Model responses were parsed using regular expression extraction. For multiple-choice datasets (MedQA and MedMCQA), the extractor identified the first standalone letter A, B, C, or D in the response. For PubMedQA, the extractor identified the first occurrence of yes, no, or maybe. Responses that did not contain a valid answer were categorized as UNKNOWN, representing instruction-following failures. All responses, including UNKNOWN outputs, were retained for consistency scoring and analysis.
@@ 3.5 Evaluation Metrics
Four complementary metrics were defined to quantify model behavior under prompt variation.
Consistency Score. For each question, the majority answer was identified, defined as the response selected most frequently across the five prompt styles, and the proportion of prompt styles that agreed with it was computed. Formally, given responses $r_{1},r_{2},r_{3},r_{4},r_{5}$ for a question, the consistency score was defined as:
@@
where $\hat{r}$ denotes the majority answer and $\mathbf{1}[\cdot]$ is the indicator function. A score of 1.0 indicates perfect consistency across all prompt styles, while a score of 0.2 indicates maximum inconsistency. For example, responses [A, B, B, B, A] yield a majority answer of B and a consistency score of $3/5=0.60$.
Overall Accuracy. The majority answer for each question is compared with the dataset ground-truth label, reflecting task-level correctness while accounting for variability across prompt styles.
@@
Statistical significance was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. All p-values are reported unadjusted. Given the exploratory nature of this analysis, no correction for multiple comparisons was applied.
Together, these metrics provide a holistic assessment of model behavior. A model may be highly consistent but consistently wrong, accurate only under specific prompt styles, or prone to instruction-following failures. Evaluating any single metric in isolation would obscure these clinically relevant failure modes.
@@ 4.1 Overall Consistency and Accuracy
Statistical significance of observed differences was assessed using the Wilcoxon signed-rank test for consistency scores and McNemar’s test for accuracy comparisons. Consistency differences between models were statistically significant on MedQA and MedMCQA (p<<0.001 for most pairwise comparisons) but not on PubMedQA, suggesting that dataset characteristics influence the degree of prompt sensitivity. Accuracy differences were generally non-significant on MedQA, whereas on PubMedQA, Llama 3.2 significantly outperformed both Phi-3 Mini and Gemma 2 (p<<0.001), indicating stronger performance on biomedical evidence-based questions. Full significance test results are provided in Table LABEL:tab:stats.
@@
MedQA Llama 3.2 Phi-3 Mini Cons. 0.0005 ***
Gemma 2 Phi-3 Mini Cons. <<.0001 ***
Mistral Phi-3 Mini Cons. <<.0001 ***
@@
Llama 3.2 Phi-3 Mini Acc. 0.9020 ns
Gemma 2 Phi-3 Mini Acc. 0.0689 ns
Mistral Phi-3 Mini Acc. 0.5383 ns
@@
MedMCQA Llama 3.2 Phi-3 Mini Cons. 0.0437 *
Gemma 2 Phi-3 Mini Cons. <<.0001 ***
Mistral Phi-3 Mini Cons. 0.0001 ***
@@
Llama 3.2 Phi-3 Mini Acc. 0.5962 ns
Gemma 2 Phi-3 Mini Acc. 0.0327 *
Mistral Phi-3 Mini Acc. 0.1060 ns
@@
PubMedQA Llama 3.2 Phi-3 Mini Cons. 0.1055 ns
Gemma 2 Phi-3 Mini Cons. 0.8497 ns
Mistral Phi-3 Mini Cons. 0.6172 ns
@@
Llama 3.2 Phi-3 Mini Acc. <<.0001 ***
Gemma 2 Phi-3 Mini Acc. 0.0002 ***
Mistral Phi-3 Mini Acc. 0.1930 ns
@@
Accuracy patterns diverged substantially from consistency rankings (Figure 2). Llama 3.2 achieved the highest overall accuracy across datasets (49.0–65.0%), followed by Phi-3 Mini (48.0–53.0%) and Mistral 7B (42.5–45.5%). Gemma 2, despite exhibiting the highest consistency scores, achieved the lowest accuracy (33.0–43.5%). This inverse relationship between consistency and accuracy in Gemma 2 highlights a critical finding: high consistency does not imply high accuracy. Models can be reliably wrong, producing the same incorrect answer across all prompt variations, which represents a particularly dangerous failure mode in clinical decision support contexts.
Meditron-7B, included as a domain-adapted pretraining model without instruction tuning, exhibited markedly different behavior from all instruction-tuned models. While it achieved moderate consistency on MedQA and MedMCQA (0.761 and 0.762, respectively), its UNKNOWN rates were substantially higher (22.8% and 22.4%), indicating frequent instruction-following failures. On PubMedQA, Meditron-7B produced near-complete failure, with a consistency score of 0.010, 2.0% accuracy, and a 99.0% UNKNOWN rate. This suggests that the model is unable to produce structured yes/no/maybe responses without instruction tuning. These results indicate that domain-specific pretraining alone is insufficient for structured clinical question answering and that instruction-following capability is a prerequisite for clinical deployment.
Due to the extremely high UNKNOWN rate observed for Meditron-7B on PubMedQA (99.0%), statistical comparisons involving this model on that dataset should be interpreted cautiously, as the near-absence of valid responses renders consistency metrics unreliable.